OpenStudy (anonymous):

Evaluate without using calculator: sin45degrees cos45degrees + tan45degrees a. (2+sqrt2)/2 b. 2 c. (1+2sqrt2)/4 d. 3/2

OpenStudy (anonymous):

\[= \frac{1}{2} + 1 = ???\]

OpenStudy (dumbcow):

i would use a unit circle to look up sin45 and cos45 sin45 = cos45 = sqrt2/2 tan45 = sin45/cos45 = 1 --> (sqrt2/2)(sqrt2/2) + 1 = 3/2

OpenStudy (anonymous):

Use them: \[\sin(45) = \cos(45) = \frac{1}{\sqrt{2}}\] \[\tan(45) = 1\]

OpenStudy (anonymous):

Thanks both of you!